please solve it manually and using ETABS.1.) Select the lightest W beam to support the following loads 4 in. concrete slab 15 ft 40 psf dead load 15 ft 100 psf live l

please solve it manually and using ETABS.
Answer #1

Weight of floor=(4/12)*150=50 psf

Additional dead load=40 psf

Total dead load=50+40=90 psf

Live load=100 psf

Factored load on floor=1.2*90+1.6*100=268 psf

Tributtary width of beam=beam spacing=15 ft

Uniform load on beam=268*15/1000=4.02 kip/ft

Span=35 ft

Maximimum moment in beam=4.02*35²/8=615.6 kip-ft

From table 3-2 of aisc, lighter w section that has flexural capacity greater than 615.6 is W24x68